### v2

@@ Appendix E Analysis for Accelerated Mirror Descent
where the last inequality is due to (12) and Lemma 17. Then we use a similar idea as in Equation 68 to bound $\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}$:
@@
where we use (96) and Lemma 3. Combine the previous results, we obtain
@@
For $e_{2}$, we also have $e_{2}=2/(2+1)\left\|\mathbf{x}_{1}-\mathbf{z}_{1}\right\|=0$ based on (12). $e_{1}=0$ is a direct consequence of our initialization.
Lemma 20.
@@
Proof E.2.
In view of Lemma 3 as well as the given condition, we have
@@
Specifying the value of $\alpha_{t}$ and $\eta_{t}$ completes the proof.
Lemma 21.
@@
Proof E.3.
We prove this lemma by induction.
@@
by choosing $\mathbf{x}=\mathbf{x}_{0}=\mathbf{z}_{0}$. Then we conclude that $f(\mathbf{x}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. By Lemma 16, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$. By definition, we have $e_{1}=0$, which completes the proof for this part.
@@
Recall the definition of $\tau$ in Theorem 9:
$\tau=\left\lceil\frac{4\sqrt{2}DL}{G}\right\rceil=\left\lceil\frac{8\sqrt{2}D}{\frac{2G}{L}}\right\rceil\geq\left\lceil\frac{8\sqrt{2}D}{2\sqrt{2}D}\right\rceil\geq 4.$ (108)
where we use $2G/L\leq 2\sqrt{2}D$ (Yu et al., 2024, Proposition A.6). Then we have
$e_{s}=\frac{2}{s+1}\left\|\mathbf{x}_{t-1}-\mathbf{z}_{t-1}\right\|\leq\frac{2}{\tau}\cdot 2\sqrt{2}D\leq\frac{4\sqrt{2}D}{\frac{4\sqrt{2}DL}{G}}\leq\frac{G}{L}.$ (109)
Case (ii). $12\leq s\leq\tau-$ 1
@@
where we use the induction basis $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G,e_{s}\leq G/L$ and Lemma 3. We also have
@@
which also holds for all $11\leq t\leq s-$ 1 by induction. Summing from $t=1$ to $t=s$, we obtain
@@
which yields $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ is immediately derived using Lemma 16. Finally, the three statements are all proven, which completes the proof.
Appendix F Analysis for Optimistic Mirror Descent
Lemma 22.
@@ Appendix F Analysis for Optimistic Mirror Descent
Similar to the proof of Lemma 26, we bound the four terms separately. To bound $A_{t}$ and $C_{t}$, we choose $\mathbf{x}=\mathbf{x}_{t-1}$ in (120), $\mathbf{x}=\mathbf{y}_{t}$ in (121). Adding them up, we have
@@
which completes the proof by plugging in the definitions of $B_{t}$ and $D_{t}$.
Lemma 23.
Under Assumptions 1, 2 and 3, consider the updates in (15) and define $G:=\sup\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\}$. If $0<\eta\leq\gamma/\ell(2G)$ for some $\gamma\in(0,1/3]$, the following statements holds for all $t\in\mathbb{N}_{+}$:
@@
Proof F.2.
The proof resembles that of Lemma 18, which involves induction. For convenience, we denote $L:=\ell(2G)$.
@@
Consider the first round of (15). Similar to the proof of Lemma 18, we argue that $f(\mathbf{y}_{0})-f^{*}=f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma 7, we have $\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}=\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}\leq G<\infty$. According to Equation 119, we have
@@
It suffices to control the inner product term. By Lemma 17 as well as our initialization $\mathbf{x}_{0}=\mathbf{y}_{0}$, we have
@@
Then we use Lemma 15 to establish the following upper bound:
@@
Now we clearly have $f(\mathbf{y}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$ as $\eta>0$. By Lemma 16, we conclude that $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Next, we prove the other half of the statements. By Equation 119, we have
@@
In the language of Equation 119, we shall establish an upper bound for the terms $\eta B_{1}$ and $\eta D_{1}$, i.e., the inner product terms in the above inequality. By Lemma 17, we have
@@
where the second inequality uses $\left\|\mathbf{y}_{1}-\mathbf{y}_{0}\right\|\leq\frac{G}{L}$ and Lemma 3. Then we invoke Lemma 15 to derive the following bounds for $\eta B_{1}$ (set $\beta=1$) as well as $\eta D_{1}$ (set $\beta=\frac{1}{2}$):
@@
where we use the initialization setup of $\mathbf{x}_{0}=\mathbf{y}_{0}$. Consequently, we conclude that $f(\mathbf{x}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$ is straightforward after a standard application of Lemma 16.
@@
where the second inequality uses Equation 157 and the third one use Lemma 3. Now we examine the rest statements for the $t$-th round of the algorithm. By Equation 119, we have
@@
To leverage the effective smoothness property, we need to carefully control the distance quantity, namely $\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|$ and $\left\|\mathbf{x}_{t}-\mathbf{y}_{t-1}\right\|$. Based on the bound of $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$, by Equations 157 and 3 we have
$\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|\leq\eta\left\|\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1})\right\|_{*}\leq\eta L\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|\leq\frac{G}{L}\sum_{s=2}^{t+1}\gamma^{s}\leq\frac{\gamma^{2}G}{(1-\gamma)L}<\frac{G}{L},$ (139)
@@
With these prepared, we invoke Lemma 15 and choose $\beta=2\gamma^{2}/(1-2\gamma)$ therein to derive
@@
where we use the value range requirement of $\gamma\in(0,1/3]$. Similarly, choosing $\beta=\gamma$ in Lemma 15 gives
@@
=$\overset{\mathbf{x}_{0}=\mathbf{y}_{0}}{=}-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}.$
@@
Obviously, we have $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$ as $\eta>0$. Applying Lemma 16 yields $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$. It remains to show the suboptimality as well as the gradient bound for $\mathbf{y}_{t}$. By Equation 119, we have
@@
Before using Lemma 15 to bound the terms $E_{t}$ and $F_{t}$, we need to verify the applicability of Lemma 3. By Lemma 17, we have
@@
Choosing $\beta=\gamma$ in Lemma 15, $E_{t}$ can be bounded as follows:
@@
Similar routine should be done for $F_{t}$. By Equations 157 and 3, we have
@@
Choosing $\beta=\gamma^{2}/(1-2\gamma)$ in Lemma 15, $F_{t}$ can be bounded as follows:
@@
We immediately see that $f(\mathbf{y}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. Then $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ is a direct result based on Lemma 16. Combining Part 1 with Part 2 finishes the proof.
Appendix G Analysis for Mirror Prox
@@
Adding $\langle\nabla f(\mathbf{x}_{1})-\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}\rangle$ to both sides of this inequality completes the proof.
Lemma 25.
Under Assumptions 1, 2 and 3, consider the updates in (17). If $0<\eta\leq 1/2\ell(2G)$ where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\right\}$, the following statements holds for all $t\in\mathbb{N}_{+}$:
@@
We consider the case where $t=1$. Similar to the proof of Lemma 18, we argue that $f(\mathbf{y}_{0})-f^{*}=f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma 7, we have $\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}=\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}\leq G<\infty$. Now we look at the sequence $\left\{\mathbf{y}_{t}\right\}_{t\in\mathbb{N}}$ that mirror prox maintains, whose dynamics are exactly like mirror descent. For convenience, we denote $\ell(2G)$ by $L$. As $\eta\leq 1/(2L)\leq 1/L$, we can apply Lemma 18 to conclude that $f(\mathbf{y}_{1})\leq f(\mathbf{x}_{0})=f(\mathbf{y}_{0})$ and $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Next, we apply Lemma 26 to the first round of the mirror prox algorithm. It’s straightforward to verify that the conditions are satisfied for the case where $s=1$. Hence, we have $f(\mathbf{x}_{1})\leq f(\mathbf{x}_{0})$. By Lemma 16, we conclude that $\left\|\nabla f(\mathbf{x}_{1})\right\|_{*}\leq G$.
@@
Suppose the statements hold for all $t\leq s-1$ where $s>2$. So we have $\left\|\nabla f(\mathbf{y}_{s-1})\right\|_{*}\leq G$ and $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$. We also have $f(\mathbf{x}_{s-1})\leq f(\mathbf{x}_{s-2})\leq f(\mathbf{x}_{0})$. Now we consider the case where $t=s$. Similar to the base case, we invoke Lemma 18 to immediately conclude that $f(\mathbf{y}_{s})\leq f(\mathbf{x}_{s-1})\leq f(\mathbf{x}_{0})$ and $\left\|\nabla f(\mathbf{y}_{s})\right\|_{*}\leq G$. Next, we prove the other half of the statements. Note that the conditions in Lemma 26 are all satisfied by the previous derivations. So we invoke Lemma 26 to deduce that $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{s-1})$, which immediately implies $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{0})$. Then $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G$ is a direct result based on Lemma 16. Combining Part 1 with Part 2 finishes the proof.
Lemma 26.
Under Assumptions 1, 2 and 3, consider the following updates for any $s\in\mathbb{N}_{+}$:
@@
$+\underbrace{\langle\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle}_{C_{s}}+\underbrace{\langle\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle}_{D_{s}}.$
@@
For $B_{s}$ and $D_{s}$, we shall use the effective smoothness property. For convenience, denote by $L:=\ell(2G)$ as we did in Lemma 3. Observe that
@@
by Lemma 17 and the given condition. Then we call Lemma 15 to bound $\eta D_{s}$ as follows:
@@
As for $B_{s}$, we still need to verify that the distance between $\mathbf{x}_{s}$ and $\mathbf{y}_{s}$ is small enough in order to apply Lemma 3. We utilize (157) to show that
@@
Then by Lemma 17 and $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$, it follows
@@
Using Lemma 3, we deduce that
@@
Utilizing Lemma 3 again along with Cauchy-Schwarz Inequality, we have

### v3

@@ Appendix F Analysis for Accelerated Mirror Descent
where the last inequality is due to (8) and Lemma E.1. Then we use a similar idea as in Equation 34 to bound $\left\|\nabla f(\mathbf{y}_{t-1})\right\|_{*}$:
@@
where we use (69) and Lemma 2.7. Combine the previous results, we obtain
@@
For $e_{2}$, we also have $e_{2}=2/(2+1)\left\|\mathbf{x}_{1}-\mathbf{z}_{1}\right\|=0$ based on (8). $e_{1}=0$ is a direct consequence of our initialization. ∎
Lemma F.2.
@@
Proof.
In view of Lemma 2.7 as well as the given condition, we have
@@
Specifying the value of $\alpha_{t}$ and $\eta_{t}$ completes the proof. ∎
Lemma F.3.
@@
  4. 4.
    (bounded trajectory) $\qquad\qquad\ \ \mathbf{x}_{t},\mathbf{z}_{t}\in\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$.
Proof.
First, we prove the bounded trajectory property. Define the Lyapunov function $E_{t}$ as
$E_{t}:=\frac{\eta t(t+1)}{4L}\left[f(\mathbf{x}_{t})-f^{*}\right]+B(\mathbf{x}_{*},\mathbf{z}_{t})\geq B(\mathbf{x}_{*},\mathbf{z}_{t})\geq\frac{1}{2}\left\|\mathbf{x}_{*}-\mathbf{z}_{t}\right\|^{2}.$ (80)
Equation 75 states that $E_{t}\leq E_{t-1}\leq\cdots E_{1}\leq E_{0}=B(\mathbf{x},\mathbf{x}_{0})$. Combining with (80), we deduce that $\left\|\mathbf{z}_{t}-\mathbf{x}_{*}\right\|\leq\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}$, implying $\mathbf{z}_{t}\in\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$. For $\mathbf{x}_{t}$, recall the update rule of $\mathbf{x}_{t}=(1-\alpha_{t})\mathbf{x}_{t-1}+\alpha_{t}\mathbf{z}_{t}$ and that $\mathbf{x}_{0}=\mathbf{z}_{0}$. We deduce that $\mathbf{x}_{t}$ is a convex combination of $\left\{\mathbf{z}_{s}\right\}_{s=0}^{t}$. By convexity of $\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$, it holds that $\mathbf{x}_{t}\in\mathcal{B}\left(\mathbf{x}_{*},\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\right)$. For conditions 1 to 3, we proceed by induction, as specified below.
@@
by choosing $\mathbf{x}=\mathbf{x}_{0}=\mathbf{z}_{0}$. Then we conclude that $f(\mathbf{x}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. By Lemma D.3, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$. By definition, we have $e_{1}=0$, which completes the proof for this part.
@@
Recall the definition of $\tau$ in Theorem 3.7:
$\tau=\left\lceil\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}L}{G}\right\rceil-1=\left\lceil\frac{8\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}{\frac{2G}{L}}\right\rceil-1\geq\left\lceil\frac{8\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}{2\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}\right\rceil-1\geq 3.$ (82)
where we use $2G/L\leq 2\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}$ (Yu et al., 2024, Proposition A.6). Then we have
$e_{s}=$ $\frac{2}{s+1}\left\|\mathbf{x}_{t-1}-\mathbf{z}_{t-1}\right\|\leq\frac{2}{\tau+1}\left(\left\|\mathbf{x}_{t-1}-\mathbf{x}_{*}\right\|+\left\|\mathbf{z}_{t-1}-\mathbf{x}_{*}\right\|\right)$ (83)
$\leq$ $\frac{2}{\tau+1}\cdot 2\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}\leq\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}}{\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}L}{G}}\leq\frac{G}{L}.$
Case (ii). $2\leq s\leq\tau-1$
@@
where we use the induction basis $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G,e_{s}\leq G/L$ and Lemma 2.7. We also have
@@
which also holds for all $1\leq t\leq s-1$ by induction. Summing from $t=1$ to $t=s$, we obtain
@@
which yields $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G<\infty$ is immediately derived using Lemma D.3. Finally, the three statements are all proven, which completes the proof. ∎
Proof of Theorem 3.7.
By Lemma F.3, we immediately conclude that
which holds for all $t\in\mathbb{N}_{+}$ and $\mathbf{x}\in\mathcal{X}$. Summing from $1$ to $T$ and choosing $\mathbf{x}=\mathbf{x}_{*}$, we obtain
$\frac{\eta T(T+1)}{4L}\left[f(\mathbf{x}_{T})-f^{*}\right]\leq B(\mathbf{x}_{*},\mathbf{z}_{0})-B(\mathbf{x}_{*},\mathbf{z}_{T})\leq B(\mathbf{x}_{*},\mathbf{x}_{0}).$ (94)
A simple rearrangement yields the desired convergence result. ∎
Appendix G Analysis for Optimistic Mirror Descent
Lemma G.1.
@@ Appendix G Analysis for Optimistic Mirror Descent
Similar to the proof of Lemma H.3, we bound the four terms separately. To bound $A_{t}$ and $C_{t}$, we choose $\mathbf{x}=\mathbf{x}_{t-1}$ in (97), $\mathbf{x}=\mathbf{y}_{t}$ in (98). Adding them up, we have
@@
which completes the proof by plugging in the definitions of $B_{t}$ and $D_{t}$. ∎
Lemma G.2.
Under Assumptions 2.5 and 3.3, consider the updates in (10) and define $G:=\sup\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\}$. If $0<\eta\leq\gamma/\ell(2G)$ for some $\gamma\in(0,1/3]$, the following statements holds for all $t\in\mathbb{N}_{+}$:
@@
Proof.
The proof resembles that of Lemma E.2, which involves induction. For convenience, we denote $L:=\ell(2G)$.
@@
Consider the first round of (10). Similar to the proof of Lemma E.2, we argue that $f(\mathbf{y}_{0})-f^{*}=f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma 3.4, we have $\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}=\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}\leq G<\infty$. According to Equation 96, we have
@@
It suffices to control the inner product term. By Lemma E.1 as well as our initialization $\mathbf{x}_{0}=\mathbf{y}_{0}$, we have
@@
Then we use Lemma D.2 to establish the following upper bound:
@@
Now we clearly have $f(\mathbf{y}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$ as $\eta>0$. By Lemma D.3, we conclude that $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Next, we prove the other half of the statements. By Equation 96, we have
@@
In the language of Equation 96, we shall establish an upper bound for the terms $\eta B_{1}$ and $\eta D_{1}$, i.e., the inner product terms in the above inequality. By Lemma E.1, we have
@@
where the second inequality uses $\left\|\mathbf{y}_{1}-\mathbf{y}_{0}\right\|\leq\frac{G}{L}$ and Lemma 2.7. Then we invoke Lemma D.2 to derive the following bounds for $\eta B_{1}$ (set $\beta=1$) as well as $\eta D_{1}$ (set $\beta=\frac{1}{2}$):
@@
where we use the initialization setup of $\mathbf{x}_{0}=\mathbf{y}_{0}$. Consequently, we conclude that $f(\mathbf{x}_{1})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$ is straightforward after a standard application of Lemma D.3.
@@
where the second inequality uses Equation 139 and the third one use Lemma 2.7. Now we examine the rest statements for the $t$-th round of the algorithm. By Equation 96, we have
@@
To leverage the effective smoothness property, we need to carefully control the distance quantity, namely $\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|$ and $\left\|\mathbf{x}_{t}-\mathbf{y}_{t-1}\right\|$. Based on the bound of $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$, by Equations 139 and 2.7 we have
$\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|\leq$ $\eta\left\|\nabla f(\mathbf{y}_{t})-\nabla f(\mathbf{y}_{t-1})\right\|_{*}\leq\eta L\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|$ (116)
$\leq$ $\frac{G}{L}\sum_{s=2}^{t+1}\gamma^{s}\leq\frac{\gamma^{2}G}{(1-\gamma)L}<\frac{G}{L},$
@@
With these prepared, we invoke Lemma D.2 and choose $\beta=2\gamma^{2}/(1-2\gamma)$ therein to derive
@@
where we use the value range requirement of $\gamma\in(0,1/3]$. Similarly, choosing $\beta=\gamma$ in Lemma D.2 gives
@@
$\overset{\mathbf{x}_{0}=\mathbf{y}_{0}}{=}-\frac{1-2\gamma}{2}\left\|\mathbf{x}_{t-1}-\mathbf{y}_{t-1}\right\|^{2}.$
@@
Obviously, we have $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$ as $\eta>0$. Applying Lemma D.3 yields $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$. It remains to show the suboptimality as well as the gradient bound for $\mathbf{y}_{t}$. By Equation 96, we have
@@
Before using Lemma D.2 to bound the terms $E_{t}$ and $F_{t}$, we need to verify the applicability of Lemma 2.7. By Lemma E.1, we have
@@
Choosing $\beta=\gamma$ in Lemma D.2, $E_{t}$ can be bounded as follows:
@@
Similar routine should be done for $F_{t}$. By Equations 139 and 2.7, we have
@@
Choosing $\beta=\gamma^{2}/(1-2\gamma)$ in Lemma D.2, $F_{t}$ can be bounded as follows:
@@
We immediately see that $f(\mathbf{y}_{t})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}$. Then $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ is a direct result based on Lemma D.3. Combining Part 1 with Part 2 finishes the proof. ∎
Proof of Equation 11.
Applying Equation 139 to the updates defined in (10), we have
By Lemma G.2, we have $\left\|\mathbf{y}_{t}-\mathbf{y}_{t-1}\right\|\leq\frac{G}{L}\sum_{s=1}^{t}1/3^{s}\leq\frac{G}{2L}<\frac{G}{L}$ as we specify $\gamma=1/3$, which satisfies the condition in Lemma D.2. Hence, the following result holds:
where we set $\beta=2/3$ in Lemma D.2. Plugging this into (130) and recalling the property of Bregman divergence, we have
By Jensen’s Inequality, we derive the following convergence rate:
$f(\overline{\mathbf{y}}_{T})-f^{*}\leq\frac{1}{\eta T}\sum_{t=1}^{T}\langle\eta\nabla f(\mathbf{y}_{t}),\mathbf{y}_{t}-\mathbf{x}_{*}\rangle\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta T},$ (134)
where $\overline{\mathbf{y}}_{T}=\sum_{t=1}^{T}\mathbf{y}_{t}/T$. ∎
Appendix H Analysis for Mirror Prox
@@
Adding $\langle\nabla f(\mathbf{x}_{1})-\mathbf{g}_{2},\mathbf{x}_{1}-\mathbf{x}\rangle$ to both sides of this inequality completes the proof. ∎
Lemma H.2.
Under Assumptions 2.5 and 3.3, consider the updates in (12). If $0<\eta\leq 1/2\ell(2G)$ where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2(f(\mathbf{x}_{0})-f^{*})\ell(2\alpha)\right\}$, the following statements holds for all $t\in\mathbb{N}_{+}$:
@@
We consider the case where $t=1$. Similar to the proof of Lemma E.2, we argue that $f(\mathbf{y}_{0})-f^{*}=f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma 3.4, we have $\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}=\left\|\nabla f(\mathbf{y}_{0})\right\|_{*}\leq G<\infty$. Now we look at the sequence $\left\{\mathbf{y}_{t}\right\}_{t\in\mathbb{N}}$ that mirror prox maintains, whose dynamics are exactly like mirror descent. For convenience, we denote $\ell(2G)$ by $L$. As $\eta\leq 1/(2L)\leq 1/L$, we can apply Lemma E.2 to conclude that $f(\mathbf{y}_{1})\leq f(\mathbf{x}_{0})=f(\mathbf{y}_{0})$ and $\left\|\nabla f(\mathbf{y}_{1})\right\|_{*}\leq G$. Next, we apply Lemma H.3 to the first round of the mirror prox algorithm. It’s straightforward to verify that the conditions are satisfied for the case where $s=1$. Hence, we have $f(\mathbf{x}_{1})\leq f(\mathbf{x}_{0})$. By Lemma D.3, we conclude that $\left\|\nabla f(\mathbf{x}_{1})\right\|_{*}\leq G$.
@@
Suppose the statements hold for all $t\leq s-1$ where $s>2$. So we have $\left\|\nabla f(\mathbf{y}_{s-1})\right\|_{*}\leq G$ and $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$. We also have $f(\mathbf{x}_{s-1})\leq f(\mathbf{x}_{s-2})\leq f(\mathbf{x}_{0})$. Now we consider the case where $t=s$. Similar to the base case, we invoke Lemma E.2 to immediately conclude that $f(\mathbf{y}_{s})\leq f(\mathbf{x}_{s-1})\leq f(\mathbf{x}_{0})$ and $\left\|\nabla f(\mathbf{y}_{s})\right\|_{*}\leq G$. Next, we prove the other half of the statements. Note that the conditions in Lemma H.3 are all satisfied by the previous derivations. So we invoke Lemma H.3 to deduce that $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{s-1})$, which immediately implies $f(\mathbf{x}_{s})\leq f(\mathbf{x}_{0})$. Then $\left\|\nabla f(\mathbf{x}_{s})\right\|_{*}\leq G$ is a direct result based on Lemma D.3. Combining Part 1 with Part 2 finishes the proof. ∎
Lemma H.3.
Under Assumptions 2.5 and 3.3, consider the following updates for any $s\in\mathbb{N}_{+}$:
@@
$+\underbrace{\langle\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle}_{C_{s}}$
$+\underbrace{\langle\nabla f(\mathbf{x}_{s})-\nabla f(\mathbf{x}_{s-1}),\mathbf{y}_{s}-\mathbf{x}_{s-1}\rangle}_{D_{s}}.$
@@
For $B_{s}$ and $D_{s}$, we shall use the effective smoothness property. For convenience, denote by $L:=\ell(2G)$ as we did in Lemma 2.7. Observe that
@@
by Lemma E.1 and the given condition. Then we call Lemma D.2 to bound $\eta D_{s}$ as follows:
@@
As for $B_{s}$, we still need to verify that the distance between $\mathbf{x}_{s}$ and $\mathbf{y}_{s}$ is small enough in order to apply Lemma 2.7. We utilize (139) to show that
@@
Then by Lemma E.1 and $\left\|\nabla f(\mathbf{x}_{s-1})\right\|_{*}\leq G$, it follows
@@
Using Lemma 2.7, we deduce that
@@
Utilizing Lemma 2.7 again along with Cauchy-Schwarz Inequality, we have